An angle is 6 degrees more then four times its complement. Whats the measure of the complement
Answer by cleomenius(959)

x + 4x + 6 = 90
5x = 84
x = 16.8 degrees, this is our first angle.
4(16.8) = 67.2 + 6 = 73.2 degrees, our second angle.
16.8 + 73.2 = 90 degrees.
Cleomenius.